OpenStudy (firejay5):

Solve each system of equations. Show your work/steps. 16. 4a + 2b - 6c = 2 6a + 3b - 9c = 3 8a + 4b - 12c = 6 18. 3x + y + z = 4 2x + 2y + 3z = 3 x + 3y + 2z = 5

OpenStudy (firejay5):

Just this one set of problems then you may go to bed

OpenStudy (ajprincess):

What is tjhe third eqn in ur question?

OpenStudy (ajprincess):

Is it 8a + 4b - 12c = 6?

OpenStudy (firejay5):

yes

OpenStudy (ajprincess):

4a + 2b - 6c = 2 6a + 3b - 9c = 3 Try eliminating the variable c from both the eqns.

OpenStudy (firejay5):

Am I done

OpenStudy (ajprincess):

No solutions for those eqns. yup u r done.

OpenStudy (firejay5):

Do #18

OpenStudy (ajprincess):

3x + y + z = 4 2x + 2y + 3z = 3 Try eliminating z from both the eqns.

OpenStudy (ajprincess):

@Firejay5

OpenStudy (firejay5):

yes

OpenStudy (ajprincess):

What do u get?

OpenStudy (firejay5):

@goformit100 Go from @ajprincess's work

OpenStudy (goformit100):

USE MATRIX METHOD FOR SOLVING THE ABOVE EQUATIONS.

OpenStudy (firejay5):

after that

OpenStudy (goformit100):

http://en.wikipedia.org/wiki/System_of_linear_equations

OpenStudy (goformit100):

OK

OpenStudy (firejay5):

can you do it @hartnn

OpenStudy (firejay5):

@goformit100 You aren't helping that much

OpenStudy (goformit100):

 SOW SOLVE IT